A die is rolled 1000 times. The total number of spots is 3680 instead of the expected 3500. Can this be explained as a chance variation, or is the die loaded?
Answer #1

here for a single roll ;

x P(x) xP(x) x2P(x)
1 0.1667 0.167 0.167
2 0.1667 0.333 0.667
3 0.1667 0.500 1.500
4 0.1667 0.667 2.667
5 0.1667 0.833 4.167
6 0.1667 1.000 6.000
total 3.500 15.167
E(x) =μ= ΣxP(x) = 3.5000
E(x2) = Σx2P(x) = 15.1667
Var(x)=σ2 = E(x2)-(E(x))2= 2.9167
std deviation=         σ= √σ2 = 1.708

from above expected value for 1000 rolls=np=1000*3.5 =3500

and standard deviation =1.708*sqrt(1000)=54.0062

as 3680 is more than 3 standard deviation from mean ; therefore this is an unusual event

and we may suspect that the die is loaded.